# Destabilisation of local magnetic anisotropy in heavy-fermion compounds

### v2

@@ Abstract
The local magnetic anisotropy of a typical crystalline compound is usually attributed to the combined effect of crystal electric fields and spin-orbit coupling. We show that this simple local picture is transformed in heavy-fermion compounds by the development of coherent electron scattering from local spin degrees of freedom. Provided the dominance of the coherence energy scale over the magnetic energy scale is strong enough, the fractionalisation and delocalisation of the spins destabilises their single-ion anisotropy by generating an opposing anisotropy in the exchange. Experimentally, this can manifest as competing splittings in the Curie-Weiss constants and effective moments. We show that the destabilisation of the anisotropy can result in either ferromagnetic or antiferromagnetic order that is perpendicular to the high-temperature easy axis. In the absence of destabilisation, we show that the order is more likely to be antiferromagnetic. In agreement with our theory – an extension of Read-Newns theory to the underscreened Kondo lattice – we observe in a wide range of actinide and lanthanide compounds that the temperature at which the anisotropy of the uniform magnetic response changes tracks the coherence energy scale. We also notice that in some compounds the destabilisation of local crystalline anisotropy can be the beginning of a cascade of decreasing energy scales. Once the anisotropy changes, further growth of magnetic correlations with cooling can lead to a rapid quenching of the hard-axis susceptibility, and at even lower temperatures, the high-temperature easy axis is revealed via a metamagnetic jump when a high magnetic field is applied along it.
@@
The rich phenomenology of magnetism in heavy-fermion metals is generally sensitive to the underlying crystal structure, crystal electric fields (CEF), spin-orbit coupling, or valency of the local-moment ions. Nonetheless, across a wide range of lanthanide and actinide compounds some robust universal features emerge Hafner et al. (2019); Aoki et al. (2019a); Ōnuki et al. (2020). One example is the reorientation of the magnetic easy axis with temperature, that is strongly correlated with the presence of the coherence maximum in the material’s resistivity. The resistivity rises logarithmically under cooling, and reaches a maximum at a characteristic temperature $T^{\ast}_{\rm coh}$, before dropping dramatically as the Kondo scattering of electrons from local spin degrees of freedom becomes coherent (see the bottom panel of Fig. 1). Many heavy-fermion compounds that order magnetically below this coherence temperature, do so at 90 degrees to the preferred direction of the local moments at high temperatures. The easy axis in the ordered state is thus at a right angle to the easy axis at temperatures above the coherence and exchange energy scales, where it is set by single-ion anisotropy and is along the direction with the highest value of the uniform susceptibility. In this Letter, we show that it is the development of coherence and flat heavy-fermion bands that destablises the local (single-ion) anisotropy and is responsible for the reorientation of the magnetic easy axis.
@@
Figure 1: Top: The susceptibility crossing temperature, $T_{\rm cross}$, against the coherence temperature $T^{*}_{\rm coh}$, for three heavy-fermion compounds. The direct proportionality fits suggest that the same energy scale is responsible for the development of coherence and easy-axis reorientation. (For UCo1-xIrxGe, the fit does not include the stars, and for EuCu2(Ge1-xSix)2 it does not include the highest $T_{\rm cross}$, which is a tentative prediction.) Bottom: A plot of the $a$-axis resistivity and modified Curie-Weiss fits of $b$- and $c$-axis susceptibilities for UCo0.98Ir0.02Ge identifying $T^{*}_{\rm coh}$ and $T_{\rm cross}$. A single-ion anisotropy, that favours antialignment of the two components of the spin-1 moment along the $c$-axis, leads to a lower paramagnetic response along this direction at high temperature. Around $T^{\star}$, when one of the components delocalises and becomes part of the heavy-fermion sea, the anisotropy interaction becomes non-local and a stronger RKKY exchange along the c-axis is generated. [A detailed description of all the plotted data and any extrapolations used can be found in Table I in the End Matter.]
@@
We have found that single-ion anisotropy is most likely to be destabilised, giving a crossing of the uniform susceptibilties and magnetic order (if present) along the high-temperature hard axis, in heavy-fermion compounds that display more itinerant rather than localised magnetism. They are characterised by clear resistivity maxima at a temperature $T^{*}$ that is significantly higher than the temperature at which potential magnetic order sets in. As shown in Table I in the End Matter, they also often possess higher Sommerfeld coefficients and smaller ordered moments, again pointing to a greater degree of itinerancy. We show that ferromagnetic order and antiferromagnetic order are both possible along the high-temperature hard axis, although the latter relies on Fermi surface nesting and is therefore more specific in nature, and perhaps less commonly seen.
The universal features of magnetism in heavy-fermion metals are not limited to the reorientation of the easy axis. In U and Eu based magnets we notice that the susceptibility crossing is often linked to the presence of a maximum in the temperature-dependence of the susceptibility along the axis that had the stronger magnetic response at high temperatures Li et al. (2021); Hovančík et al. (2022); Pospíšil et al. (2017); Iha et al. (2018); Szlawska et al. (2023); Hiranaka et al. (2013). This so-called hard-axis maximum and the following decay of the hard-axis susceptibility with cooling are in turn connected to a metamagnetic transition when a high magnetic field is applied along the axis at lower temperatures Li et al. (2021); Hovančík et al. (2022); Pospíšil et al. (2017); Hiranaka et al. (2013). In fact, looking at all the heavy-fermion metals that we have surveyed (see Table I in the End Matter), it appears that a metamagnetic transition in a high field applied along the hard axis implies a maximum in the susceptibility along this axis at a higher temperature, which in turn implies that it is the easy axis at an even higher temperature.
In some compounds, e.g. URhGe, coherence develops at a temperature $T^{\ast}_{\rm coh}$ that is much higher than the temperature at which fairly large moments order, pointing to the underscreened nature of the underlying Kondo lattice, where in spite of the Kondo energy scale $T_{K}$ being much higher than the RKKY energy scale $T_{\rm RKKY}$, there are not enough conduction electrons to fractionalise the local spin degrees of freedom entirely and a so-called residual moment remains. Motivated by the above observation, we present a generalisation of the standard large-$N$ Read-Newns theory to the underscreened Kondo lattice. We will focus on the underscreened $S=1$ Kondo lattice.
@@
$+\sum_{\eta i}D^{\eta}\left(S^{\eta}(\mathbf{r}_{i})\right)^{2}{\color[rgb]{0,0,0}-\sum_{i}\mathbf{h}_{i}\cdot\left(\mathbf{S}(\mathbf{r}_{i})+\mathbf{s}(\mathbf{r}_{i})\right)},$
where $J_{K}>0$ is the Kondo coupling between local spin $\mathbf{S}_{i}$ and the conduction electron spin $\mathbf{s}_{i}$, $\sigma$ denote the up/down spin indices, $c_{\sigma}(\mathbf{k})$ are the usual conduction electron annihilation operators, and $\mathbf{h}$ is an external magnetic field. For simplicity, we have assumed that electrons and spins occupy the same Bravais lattice sites, and $\epsilon_{\mathbf{k}}$ is the dispersion of the single conduction band with bandwidth $W$. Unless stated otherwise, we will be using a constant density of states $\rho=1/W$. $D^{x,y,z}$ is a general single-ion anisotropy for spin-1, and we have set its average value to zero. We will work away from the insulating limit, and unless specified otherwise, just below half-filling.
@@
where $a=1,2$ index the two spin-1/2 fermions, $\sigma,\sigma^{\prime}$ denote the up/down spin indices, $\sigma^{\eta}$ are Pauli matrices with $\eta\in\{x,y,z\}$, and $f_{a\sigma}(\mathbf{r}_{i})$ obey the usual fermionic anticommutation relations. We will also promote the $SU(2)$ spin degrees of freedom to $SU(N)$. This can be done in a number of ways and we will choose to have $N/2$ replicas of spin-up states and an equal number of spin-down states on each lattice site, with half of them occupied, so that we are connected most straightforwardly to the physical $N=2$ limit. $N$ will be the control parameter for the large-$N$ theory. The Hamiltonian can be written (up to a constant) as
@@
${\color[rgb]{0,0,0}-\sum_{i\eta\alpha}h_{i}^{\eta}\left(\sum_{a}S^{\eta}_{a\alpha}(\mathbf{r}_{i})+s^{\eta}_{\alpha}(\mathbf{r}_{i})\right)},$
@@
and $\alpha,\beta$ index the spin up/down replicas. We have rescaled the Kondo coupling $J_{K}\rightarrow(2J_{K}/N)$ so that energy density remains finite as $N\rightarrow\infty$, and, up to a constant, $H_{N=2}=H$. The enlargement of the Hilbert space through the Schwinger fermion representation comes with several constraints and associated gauge symmetries:
@@
The isotropic ($D^{\eta}=0$) system is controlled by the fixed-gauge saddle point as $N\rightarrow\infty$. Read-Newns large-$N$ theory introduces the hybridisation field
$\begin{pmatrix}\hat{V}_{1i}\\ \hat{V}_{2i}\end{pmatrix}=\frac{1}{N}\sum_{\alpha\sigma}\begin{pmatrix}f^{\dagger}_{1\alpha\sigma}(\mathbf{r}_{i})c_{\alpha\sigma}(\mathbf{r}_{i})\\ f^{\dagger}_{2\alpha\sigma}(\mathbf{r}_{i})c_{\alpha\sigma}(\mathbf{r}_{i})\end{pmatrix},$ (5)
and the $U(2)$ gauge symmetry can be used to set $\hat{V}_{2i}=0$ and $\hat{V}_{1i}>0$. $f_{2}$-moments thus decouple from conduction electrons in the hybridisation channel entirely and remain unfractionalised, whereas the $f_{1}$-moments fractionalise and lead to the enlargement of the Fermi surface and flat heavy-fermion bands. (If saddle point corrections are included there is a weak residual interaction that anyway decays logarithmically with temperature Scott and Kwasigroch .)
Residual local moments. – This would be the complete story for a single impurity, but for a Kondo lattice, the macroscopic degeneracy of the $f_{2}$-moment subspace can be broken by a coupling that leaves a vanishing imprint on the $f_{1}$-$c$ heavy-fermion subspace. Because of the $U(2)$ gauge symmetry, $f_{2}$-fermions can be decoupled from the conduction electrons in the hybridisation channel but not the magnetic channel since magnetisation, $S^{\eta}(\mathbf{r}_{i})$, is a gauge-invariant physical operator. The fixed-gauge $N\rightarrow\infty$ saddle point does not contain any interactions in this channel. However, the fact that for any finite $N$, as $J_{K}/W\rightarrow 0$, the conventional RKKY ground state becomes favourable over the heavy-fermion phase can be picked up variationally, and this is indeed the original Doniach argument Doniach (1977). We therefore consider the following variational ansatz, which allows for interactions in both magnetic and hybridisation channels, for the state of the system at $T\ll T_{K}$
@@
where $\nu$ indexes two heavy fermion bands and $|\psi_{2i}\rangle$ is the state of one of the $f_{2}$-moments. By employing the usual static approximation for the $f_{2}$-moments, calculating $\langle\Psi|H_{N}|\Psi\rangle$ with $\{u^{\nu}_{\mathbf{k}\sigma},v^{\nu}_{\mathbf{k}\sigma}\}$ chosen to minimise the expectation value subject to global constraints on the number of conduction and $f_{1}$ fermions, we can derive the effective RKKY interaction between $f_{2}$-moments that is mediated by the heavy fermions. Expanding to quadratic order in $1/N$ and $D^{\eta}$ and linear order in an external magnetic field $h$, we obtain
$H_{\rm eff}\!=\!-\frac{1}{2}{\color[rgb]{0,0,0}\sum_{i\neq j,\eta}\mathcal{J}^{\eta}(\mathbf{r}_{i}-\mathbf{r}_{j})S^{\eta}_{2}(\mathbf{r}_{i})S^{\eta}_{2}(\mathbf{r}_{j})}\!-\!\sum_{\mathbf{q}\eta}\tilde{h}^{\eta}(\mathbf{q})S^{\eta}_{2}(\mathbf{q}),$ (7)
@@
where $\chi_{cc}(\mathbf{q})$ ($\chi_{cf}(\mathbf{q})$) gives the magnetic susceptibility of the conduction electrons ($f_{1}$ local moments) when a $\mathbf{q}$-modulated magnetic field is applied isothermally to conduction electrons only. Similarly, $\chi_{ff}(\mathbf{q})$ gives the magnetic susceptibility of the $f_{1}$ local moments, when the applied field only couples to them. $E^{\nu}_{\mathbf{k}}$ is the energy dispersion of the heavy-fermion quasiparticles in band $\nu$ and $n(E^{\nu}_{\mathbf{k}})$ the corresponding Fermi occupation number. (Expressions for $u^{\nu}_{\mathbf{k}},v^{\nu}_{\mathbf{k}}$ and $E^{\nu}_{\mathbf{k}}$ and and an outline of the derivation of the effective Hamiltonian can be found in the End Matter.)
The effective Hamiltonian in Eq. 7 is the main result of our work and in the isotropic and insulating limit agrees with the Hamiltonian derived in Ref. Ko et al. (2013) via a slightly different approach. There, the difference between $H_{N=2}$ and the mean-field Hamiltonian (exact for $N\rightarrow\infty$) was considered as the residual perturbative interaction and the effective Hamiltonian for the $f_{2}$-subspace was obtained. The results of that approach agreed well with DMRG calculations for a 1D Kondo lattice problem.
@@
Figure 2: The ratios $T_{\rm N/C}/T_{\rm cross}$ and $T_{\rm cross}/T^{*}_{\rm coh}$ plotted for several compounds, where $T_{\rm cross}$ is the uniform susceptibility crossing temperature, $T^{*}_{\rm coh}$ the coherence temperature, and $T_{\rm N/C}$ the Neel (circles) or Curie (squares) temperature depending on the type of magnetic order. Triangles denote lack of magnetic order. UCo1-xIrxGe transitions between the strong-coupling and weak-coupling regimes with $x$. Inset: the ratios $T_{\rm RKKY}/T_{\rm cross}$ and $(N/2)T_{\rm cross}/T_{K}$ plotted against the Kondo coupling strength, $J_{K}/W$. There are two regimes: Weak coupling at $J_{K}/W\lesssim 1/2$, where the Curie-Weiss constant is negative and $T_{\rm cross}$ approaches $T_{\rm RKKY}$ before vanishing entirely around $J_{K}/W\sim 1/\ln{N}$ , and strong coupling at $J_{K}/W\gtrsim 1/2$, where the Curie-Weiss constant is positive and $T_{\rm RKKY}\ll T_{\rm cross}\approx T_{K}/N$.
@@
Easy-axis reorientation.Clearly at temperatures much greater than $T_{K}$ and $T_{\rm RKKY}$, where the spin-1 moments are unhybridised and RKKY interactions can be neglected, a hard-axis anisotropy $D^{\eta}>0$ along $\eta$ leads to a smaller magnetic response along this axis. Let us now use the effective Hamiltonian and consider a magnetic modulation with wavevector $\mathbf{q}$ at temperatures below $T_{K}$. Assuming for now (see later discussion) that $\chi(\mathbf{q})$ and $\chi(\mathbf{q})-\bar{\chi}$ are all positive for $T\ll T_{K}$ (the bar denotes an average of the respective function over the Brillouin zone), the anisotropy of the effective field, $\tilde{h}^{\eta}(\mathbf{q})$, and the anisotropy of the RKKY exchange $\mathcal{J}^{\eta}(\mathbf{q})-\bar{\mathcal{J}}^{\eta}$ are of opposite sign to first order in $D^{\eta}$. For example, a hard-axis anisotropy $D^{\eta}>0$ along $\eta$ leads to a reduction of the residual effective moment but a reinforcement of RKKY exchange along this direction. We can understand this by looking at how the residual moment polarises the surrounding heavy Fermi sea. The greater the $D^{\eta}$ the stronger the anti-alignment of the heavy Fermi liquid induced by the anisotropy interaction between the $f_{2}$, and the now delocalised, $f_{1}$-moments (see the bottom panel of Fig. 1). Locally this gives a lower effective moment along the direction with higher $D^{\eta}$, and therefore lower paramagnetic response at high temperature, but globally, the long-range RKKY forces are stronger along this direction. The collective fluctuations generated by RKKY dominate at low temperatures and the higher the $D^{\eta}$ the stronger the magnetic response. In particular, to first order in $D^{\eta}$, a magnetically ordered ground state (at some wavevector $\mathbf{q}$ that maximises $\chi_{cc}(\mathbf{q}$)) would have moments along the direction with the highest $D^{\eta}$, and hence strongest RKKY exchange but smallest susceptibility (at the same $\mathbf{q}$) in the high-temperature paramagnetic regime because of the lowest effective moment (or indeed in the unhybridised-moment regime above the coherence temperature). This is the hard-direction ordering discussed in Ref. Kwasigroch et al. (2022).
@@
to quadratic order in $1/N$ and first non-vanishing order in $D^{\eta}$. We have used the Weiss mean-field approximation for the spins. We should mention here that spin Hamiltonians with opposing anisotropies in the effective moments (or Landé g-factors) and Curie-Weiss constants have been postulated previously, showing excellent agreement with experimental data, see for example Ref. Andrade et al. (2014). However, we believe that this is the first time that such a spin Hamiltonian is derived microscopically from a generic Kondo lattice model.
We identify two regimes as depicted in Fig. 2. In the strong-coupling limit $J_{K}/W\gtrsim 1/2$, we have to first non-vanishing order in $D^{\eta}$ and $1/N$
@@
where $T_{K}$ is the Kondo energy scale  [^1] We take it to be equal to the difference between the energy of the $N\rightarrow\infty$ ground state and the $N\rightarrow\infty,J_{K}=0$ Fermi gas state.. In the weak-coupling limit, for any fixed $N$, as $J_{K}$ is reduced, the crossing temperature approaches the RKKY energy scale, which we define as $T_{\rm RKKY}=|\Theta_{\rm CW}|=(J_{K}/N)^{2}|\chi_{cc}(\mathbf{0})-\bar{\chi}_{cc}|$, before vanishing entirely. We also find that the sign of the Curie-Weiss constant $\Theta_{\rm CW}$ changes from positive to negative around $J/W\sim 1/2$ as we transition from the strong to weak coupling regime, suggesting that the ground state is ferromagnetic at strong-coupling and antiferromagnetic at weak-coupling. The fact that we see this for a constant density of states approximation, where there is no specific underlying lattice, shows that this is a universal phenomenon, which we indeed observe for other lattices (see our later discussion). Fig. 2 illustrates the two regimes for both theory and for experimentally measured values for UCo1-xIrxGe, which spans the range from strongly coupled ($T_{\rm N/C}/T_{\rm cross}\sim 0$) to weakly coupled ($T_{\rm N/C}/T_{\rm cross}\sim 1$) as a function of the doping $x$, where $T_{\rm N/C}$ is the Neel or Curie temperature depending on the type of magnetic order.
If we extrapolate the strong-coupling large-$N$ result to the physical case $N=2$, we find that the crossing temperature is proportional to the strength of the Kondo interaction, i.e., we should have $T_{\rm cross}\propto T^{*}_{\rm coh}$, which indeed shows excellent agreement with experimental data collected across a wide range of heavy-fermion compounds as shown in Fig. 1, where the Kondo coherence temperature is varied with pressure or doping for three heavy fermion compounds and is plotted against the susceptibility crossing temperature, $T_{\rm cross}$.
@@
Magnetic ground states – We apply our effective Hamiltonian to the representative case of the simple cubic lattice with nearest neighbour hopping (we have observed similar behaviour for a generic triclinic lattice) and determine its classical $T=0$ phases in the weak anisotropy limit $D^{\eta}\rightarrow 0$ for a range of conduction electron lattice fillings $10.75<n_{c}<$ 1 and coupling constants $30.1<J_{K}/W<$ 3. To simplify the discussion of easy-axis reorientation, we restrict ourselves to simple collinear phases, where the location of the maximum of $\chi_{cc}(\mathbf{q})$ determines the magnetic ordering wavevector $\mathbf{q}$ and if $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}$ is positive (negative) at this wavevector the magnetic order is along the high-temperature hard (easy) axis.

### v3

@@ Abstract
The local magnetic anisotropy of a typical crystalline compound is usually attributed to the combined effect of crystal electric fields and spin-orbit coupling. We show that this simple local picture is transformed in heavy-fermion compounds by the development of coherent electron scattering from local spin degrees of freedom. Provided the dominance of the coherence energy scale over the magnetic energy scale is strong enough, the fractionalisation and delocalisation of the spins destabilises their single-ion anisotropy by generating an opposing anisotropy in the exchange. Experimentally, this can manifest as competing splittings in the Curie-Weiss constants and effective moments. We show that in the presence of orthorhombic or tetragonal symmetry the destabilisation of the anisotropy can result in either ferromagnetic or antiferromagnetic order that is perpendicular to the high-temperature easy axis. In the absence of destabilisation, we show that the order is more likely to be antiferromagnetic. In agreement with our theory, we also observe that the temperature at which the anisotropy of the uniform magnetic response changes tracks the coherence energy scale in a wide range of actinide and lanthanide compounds.
@@
The rich phenomenology of magnetism in heavy-fermion metals is generally sensitive to the underlying crystal structure, crystal electric fields (CEF), spin-orbit coupling, or valency of the local-moment ions. Nonetheless, across a wide range of lanthanide and actinide compounds some robust universal features emerge Hafner et al. (2019); Aoki et al. (2019a); Ōnuki et al. (2020). One example is the reorientation of the magnetic easy axis with temperature, that is strongly correlated with the presence of the coherence maximum in the material’s resistivity. The resistivity rises logarithmically under cooling, and reaches a maximum at a characteristic temperature $T^{\ast}_{\rm coh}$, before dropping dramatically as the Kondo scattering of electrons from local spin degrees of freedom becomes coherent (see the bottom panel of Fig. 1). Many heavy-fermion compounds with orthorhombic or tetragonal symmetry that order magnetically below this coherence temperature, do so at 90 degrees to the preferred direction of the local moments at high temperatures. The easy axis in the ordered state is thus perpendicular to the easy axis at temperatures above the coherence and exchange energy scales, where it is set by the single-ion anisotropy. In this Letter, we show that it is the development of coherence and flat heavy-fermion bands that destablises the local (single-ion) anisotropy and is responsible for the reorientation of the magnetic easy axis.
@@
Figure 1: Top: The susceptibility crossing temperature, $T_{\rm cross}$, against the coherence temperature $T^{*}_{\rm coh}$, for three heavy-fermion compounds. The direct proportionality fits suggest that the same energy scale is responsible for the development of coherence and easy-axis reorientation. (For UCo1-xIrxGe, the fit does not include the stars, and for EuCu2(Ge1-xSix)2 it does not include the highest $T_{\rm cross}$, which is a tentative prediction.) Bottom: A plot of the $a$-axis resistivity and modified Curie-Weiss fits of $b$- and $c$-axis susceptibilities for UCo0.98Ir0.02Ge identifying $T^{*}_{\rm coh}$ and $T_{\rm cross}$. A single-ion anisotropy, that favours antialignment of the two components of the spin-1 moment along the $c$-axis, leads to a lower paramagnetic response along this direction at high temperature. Around $T^{*}_{\rm coh}$, when one of the components delocalises and becomes part of the heavy-fermion sea, the anisotropy interaction becomes non-local and a stronger RKKY exchange along the c-axis is generated. [A detailed description of all the plotted data and any extrapolations used can be found in Table I in the End Matter.]
@@
We find that single-ion anisotropy is most likely to be destabilised in heavy-fermion compounds that display more itinerant rather than localised magnetism. They are characterised by clear resistivity maxima at a temperature $T^{*}_{\rm coh}$ that is significantly higher than the temperature at which potential magnetic order sets in. As shown in Table I in the End Matter, they also often possess higher Sommerfeld coefficients and smaller ordered moments, again pointing to a greater degree of itinerancy. We show that ferromagnetic order and antiferromagnetic order are both possible along the high-temperature hard axis, although the latter relies on Fermi surface nesting and is therefore more specific in nature, and perhaps less commonly seen. In the absence of destabilisation, antiferromagnetic order appears to be more common.
[…]
@@
$+\sum_{\eta i}D^{\eta}\left(S^{\eta}(\mathbf{r}_{i})\right)^{2}{\color[rgb]{0,0,1}-\sum_{i}\mathbf{h}_{i}\cdot\left(\mathbf{S}(\mathbf{r}_{i})+\mathbf{s}(\mathbf{r}_{i})\right)},$
where $J_{K}>0$ is the Kondo coupling between local spin $\mathbf{S}_{i}$ and the conduction electron spin $\mathbf{s}_{i}$, $\sigma$ denote the up/down spin indices, $c_{\sigma}(\mathbf{k})$ are the usual conduction electron annihilation operators, and $\mathbf{h}$ is an external magnetic field. For simplicity, we have assumed that electrons and spins occupy the same Bravais lattice sites, and $\epsilon_{\mathbf{k}}$ is the dispersion of the single conduction band with bandwidth $W$. Unless stated otherwise, we will be using a constant density of states $\rho=1/W$. $D^{x,y,z}$ is a general orthorhombic single-ion anisotropy for spin-1, and we have set its average value to zero. We will work away from the insulating limit, and unless specified otherwise, just below half-filling.
@@
[…]
@@
${\color[rgb]{0,0,1}-\sum_{i\eta\alpha}h_{i}^{\eta}\left(\sum_{a}S^{\eta}_{a\alpha}(\mathbf{r}_{i})+s^{\eta}_{\alpha}(\mathbf{r}_{i})\right)},$
@@
and $\alpha,\beta$ index the spin up/down replicas. The enlargement of the Hilbert space through the Schwinger fermion representation comes with several constraints and associated gauge symmetries:
@@
[…]
$\begin{pmatrix}\hat{V}_{1i}\\ \hat{V}_{2i}\end{pmatrix}=\frac{1}{N}\sum_{\alpha\sigma}\begin{pmatrix}f^{\dagger}_{1\alpha\sigma}(\mathbf{r}_{i})c_{\alpha\sigma}(\mathbf{r}_{i})\\ f^{\dagger}_{2\alpha\sigma}(\mathbf{r}_{i})c_{\alpha\sigma}(\mathbf{r}_{i})\end{pmatrix}.$ (5)
In the saddle-point approximation, the local $U(2)$ gauge symmetry can be used to set $\langle\hat{V}_{2i}\rangle=0$ and $\langle\hat{V}_{1i}\rangle>0$. $f_{2}$-moments thus decouple from conduction electrons at the level of large-N mean-field theory Ko et al. (2013).
[…]
@@
[…]
$H_{\rm eff}\!=\!-\frac{1}{2}{\color[rgb]{0,0,1}\sum_{i\neq j,\eta}\mathcal{J}^{\eta}(\mathbf{r}_{i}-\mathbf{r}_{j})S^{\eta}_{2}(\mathbf{r}_{i})S^{\eta}_{2}(\mathbf{r}_{j})}\!-\!\sum_{\mathbf{q}\eta}\tilde{h}^{\eta}(\mathbf{q})S^{\eta}_{2}(\mathbf{q}),$ (7)
@@
where $\chi_{cc}(\mathbf{q})$ ($\chi_{cf}(\mathbf{q})$) gives the magnetic susceptibility of the conduction electrons ($f_{1}$ local moments) when a $\mathbf{q}$-modulated magnetic field is applied isothermally to conduction electrons only. Similarly, $\chi_{ff}(\mathbf{q})$ gives the magnetic susceptibility of the $f_{1}$ local moments, when the applied field only couples to them. $E^{\nu}_{\mathbf{k}}$ is the energy dispersion of the heavy-fermion quasiparticles in band $\nu$ and $n(E^{\nu}_{\mathbf{k}})$ the corresponding Fermi occupation number. (Expressions for $u^{\nu}_{\mathbf{k}},v^{\nu}_{\mathbf{k}}$ and $E^{\nu}_{\mathbf{k}}$ and an outline of the derivation of the effective Hamiltonian can be found in the End Matter.) In the isotropic limit, our effective Hamiltonian agrees with the one derived in Ref. Ko et al. (2013) for an underscreened spin1-1- 1 Kondo chain.
@@
Figure 2: The ratios $T_{\rm N/C}/T_{\rm cross}$ and $T_{\rm cross}/T^{*}_{\rm coh}$ plotted for UCo1-xIrxGe, where $T_{\rm cross}$ is the uniform susceptibility crossing temperature, $T^{*}_{\rm coh}$ the coherence temperature, and $T_{\rm N/C}$ the Neel (circles) or Curie (squares) temperature depending on the type of magnetic order. Triangles denote lack of magnetic order. Inset: the ratios $T_{\rm RKKY}/T_{\rm cross}$ and $(N/2)T_{\rm cross}/T_{K}$ plotted against the Kondo coupling strength, $J_{K}/W$. There are two regimes: weak coupling at $J_{K}/W\lesssim 1/2$, where the Curie-Weiss constant is negative and $T_{\rm cross}$ approaches $T_{\rm RKKY}$ before vanishing entirely around $J_{K}/W\sim 1/\ln{N}$, and strong coupling at $J_{K}/W\gtrsim 1/2$, where the Curie-Weiss constant is positive and $T_{\rm RKKY}\ll T_{\rm cross}\approx T_{K}/N$.
@@
Easy-axis reorientation.Clearly at temperatures much greater than $T_{K}$ and $T_{\rm RKKY}$, where the spin-1 moments are unhybridised and RKKY interactions can be neglected, a hard-axis anisotropy $D^{\eta}>0$ along $\eta$ leads to a smaller magnetic response along this axis. Let us now consider a magnetic modulation with wavevector $\mathbf{q}$ at temperatures below $T_{K}$. Assuming for now (see later discussion) that $\chi(\mathbf{q})$ and $\chi(\mathbf{q})-\bar{\chi}$ are all positive for $T\ll T_{K}$ (the bar denotes an average of the respective function over the Brillouin zone), the anisotropy of the effective field, $\tilde{h}^{\eta}(\mathbf{q})$, and the anisotropy of the RKKY exchange $\mathcal{J}^{\eta}(\mathbf{q})-\bar{\mathcal{J}}^{\eta}$ are of opposite sign to first order in $D^{\eta}$. For example, a hard-axis anisotropy $D^{\eta}>0$ along $\eta$ leads to a reduction of the residual effective moment but a reinforcement of RKKY exchange along this direction. We can understand this by looking at how the residual moment polarises the surrounding heavy Fermi sea. The greater the $D^{\eta}$ the stronger the local anti-alignment induced by the anisotropy interaction between the $f_{2}$, and the now delocalised, $f_{1}$-moments (see the bottom panel of Fig. 1). Locally this gives a lower effective moment along the direction with higher $D^{\eta}$, and therefore lower paramagnetic response at high temperature, but globally, the long-range RKKY forces are stronger along this direction. The collective fluctuations generated by RKKY dominate at low temperatures and the higher the $D^{\eta}$ the stronger the magnetic response. In particular, a magnetically ordered collinear ground state would have moments along the direction with the highest $D^{\eta}$.
@@
in the Weiss mean-field approximation, and to quadratic order in $1/N$ and first non-vanishing order in $D^{\eta}$.
[…]
@@
where $T_{K}$ is the Kondo energy scale  [^1] We take it to be equal to the difference between the energy of the $N\rightarrow\infty$ ground state and the $N\rightarrow\infty,J_{K}=0$ Fermi gas state.. Extrapolating to the physical case $N=2$, we find that the crossing temperature is proportional to the strength of the Kondo interaction, i.e., we should have $T_{\rm cross}\propto T^{*}_{\rm coh}$, which indeed shows excellent agreement with experimental data in a wide range of heavy-fermion compounds as shown in Fig. 1.
In the weak-coupling limit, for any fixed $N$, as $J_{K}$ is reduced, the crossing temperature approaches the RKKY energy scale, which we define as $T_{\rm RKKY}=|\Theta_{\rm CW}|=(J_{K}/N)^{2}|\chi_{cc}(\mathbf{0})-\bar{\chi}_{cc}|$, before vanishing entirely. We also find that the sign of the $D^{\eta}=0$ Curie-Weiss constant $\Theta_{\rm CW}$ changes from positive to negative around $J/W\sim 1/2$ as we transition from the strong to weak coupling regime, suggesting that the ground state is ferromagnetic at strong-coupling and antiferromagnetic at weak-coupling. The fact that we see this for a constant density of states approximation, where there is no specific underlying lattice, shows that this is a universal phenomenon, which we indeed observe for other lattices (see our later discussion). Fig. 2 illustrates the two regimes for both theory and for experimentally measured values for UCo1-xIrxGe, which spans the range from strongly to weakly coupled as a function of the doping $x$.
@@
Magnetic ground states – We apply our effective Hamiltonian to the representative case of the simple cubic lattice with nearest neighbour hopping (similar behaviour observed for other orthorhombic lattices) and determine its classical $T=0$ phases in the weak anisotropy limit $D^{\eta}\rightarrow 0$ for a range of conduction electron lattice fillings $0.75<n_{c}<1$ and coupling constants $1/6<J_{K}/W<3$. To simplify the discussion of easy-axis reorientation, we restrict ourselves to simple collinear phases, where the location of the maximum of $\chi_{cc}(\mathbf{q})$ determines the magnetic ordering wavevector $\mathbf{q}$ and if $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}$ is positive (negative) at this wavevector the magnetic order is along the high-temperature hard (easy) axis.